Jr is 4 years older than kyle. Two years ago, he was 2 times as old as kyle. What are their present ages?
Answer by ikleyn(52776)

.
Jr is 4 years older than kyle. Two years ago, he was 2 times as old as kyle. What are their present ages?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

j = k + 4,        (1)
j-2 = 2*(k-2),    (2)

where j = Jr's present age, k = Kyle's present age.

Simplify (2):

j-2 = 2k - 4.

Now substitute j = k+4 from (1) into (2). You will get

(k+4) - 2 = 2k - 4.

Simplify:

k + 2 = 2k - 4,
6 = 2k - k,
k = 6,

Thus Kyle's present age is 6 yers.
Then Jr's present age is j = k+4 = 6 + 4 = 10, from (1).